The boiling point of the product( tert-butyl ethyl fumarate) is 105C-107C at 12mm Hg. What is the boiling poing at atmospheric pressure (760mm)? why?
Answer #1

Using clausius - Clapeyron equation, we can calculate boiling point at 760 mm pressure. la = DHvar (tx- ) -(1) From data, Pi0026 -4,147 4.56x102 fo = 0.909 x 10 = 0.0026 +0.909 x103 = 2.6x103 + 0.909 #10 3.51 x103 3 0.285 x 10 --3 = 3.51 X103 = 285